I'm doing functions and need some help. the book I am using is 12th edition Algebra and Trig with analytical geometry by Swokowski and Cole. the homework is page 190-192 #3-45 by 3's.
-thanks
if f(x)= sqrt(x-4-3x), find f(4), f(8), and f(13)

Answer by Alan3354(69443)

if f(x)= sqrt(x-4-3x), find f(4), f(8), and f(13)
---------------
Sub the values for x:
f(x) = sqrt(-4-2x)
f(4) = sqrt(-4-2*4) = sqrt(-12) = 2sqrt(-3)
-------------
f(8) = sqrt(-4-2*8) = sqrt(-20) etc